how to take an average between 73 and 83?
Answer by stanbon(75887)

how to take an average between 73 and 83?
-----
average = (sum of the numbers)/(number of numbers)
----
Your Problem::
(73+83)/2 = 156/2 = 78
---------------
cheers,
Stan H.
--------------